# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 2.1 Assumptions
In this subsection, we provide the assumptions. But before that we need the following notation. We use the following notation in the paper regarding vector and matrix norms. Let $\boldsymbol{v}$ be a generic $n\times 1$ vector and $\boldsymbol{H}$ to be a generic $n\times n$ matrix. Then $\|\boldsymbol{v}\|_{1},\|\boldsymbol{v}\|_{2},\|\boldsymbol{v}\|_{\infty}$ are the $l_{1}$, Euclidean, and sup norms of the vector $\boldsymbol{v}$ respectively. Next, $\|\boldsymbol{H}\|_{l_{1}},\|\boldsymbol{H}\|_{l_{\infty}}$ are the maximum column sum matrix norm, and maximum row sum matrix norm respectively. This means $\|\boldsymbol{H}\|_{l_{1}}\mathrel{\mathop{:}}=\max_{1\leq j\leq n}\sum_{i=1}^{n}|H_{i,j}|$, and $\|\boldsymbol{H}\|_{l_{\infty}}\mathrel{\mathop{:}}=\max_{1\leq i\leq n}\sum_{j=1}^{n}|H_{i,j}|$, where $H_{i,j}$ represents $(i,j)$ th element of matrix $H$. Let $\|\boldsymbol{H}\|_{l_{2}}$ be the spectral norm of matrix H, and $\|\boldsymbol{H}\|_{\infty}\mathrel{\mathop{:}}=\max_{1\leq i\leq n}\max_{1\leq j\leq n}|H_{i,j}|$.
First, $\boldsymbol{u}_{-j,t}$ is the $(p-1)\times 1$ vector of errors in $t$ th time period, except the $j$th term in $\boldsymbol{u}_{t}$. Then define $\eta_{j,t}\mathrel{\mathop{:}}=u_{j,t}-\boldsymbol{u}_{-j,t}^{\prime}\boldsymbol{\gamma}_{j}$. Also define $Eigmin(\boldsymbol{A})$ to be the minimum eigenvalue of the matrix ${\boldsymbol{A}}$.
Assumption 1.
@@
(i). $\{\boldsymbol{u}_{t}\}_{t=1}^{T},\{\boldsymbol{f}_{t}\}_{t=1}^{T}$ are sequences of (strictly) stationary and ergodic random vectors . Furthermore, $\{\boldsymbol{u}_{t}\}_{t=1}^{T},\{\boldsymbol{f}_{t}\}_{t=1}^{T}$ are independent. $\boldsymbol{u}_{t}$ is a ($p\times 1$) zero mean random vector with covariance matrix $\boldsymbol{\Sigma}_{u}$ ($p\times p$). $\textnormal{Eigmin}(\boldsymbol{\Sigma}_{u})\geq c>0$, with $c$ a positive constant, and $\max_{1\leq j\leq p}\mathbb{E}\left[u_{j,t}^{2}\right]\leq C<\infty$. (ii). For the strong mixing vector of random variables $\boldsymbol{f}_{t},\boldsymbol{u}_{t}$: $\alpha(t)\leq\exp(-Ct^{r_{0}})$, for a positive constant $r_{0}>0$.
Assumption 2.
@@
There exists positive constants $r_{1}$, $r_{2}$, $r_{3}>0$ and another set of positive constants $c_{1}$, $c_{2}$, $c_{3}$, $s_{1}$, $s_{2}$, $s_{3}>0$, and for $t=1,\cdots,T$, and $j=1,\cdots,p$, with $k=1,\cdots,K$
(i).
$\mathbb{P}\left[|u_{j,t}|>s_{1}\right]\leq\exp[-(s_{1}/c_{1})^{r_{1}}].$
(ii).
$\mathbb{P}\left[|\eta_{j,t}|>s_{2}\right]\leq\exp[-(s_{2}/c_{2})^{r_{2}}].$
(iii).
$\mathbb{P}\left[|f_{k,t}|>s_{3}\right]\leq\exp[-(s_{3}/c_{3})^{r_{3}}].$
(iv). There exists $10<\phi_{1}<$ 1 such that $\phi_{1}^{-1}=3r_{1}^{-1}+r_{0}^{-1}$, and we also assume $13r_{2}^{-1}+r_{0}^{-1}>$ 1, and $13r_{3}^{-1}+r_{0}^{-1}>$ 1.
Define $\phi_{2}^{-1}\mathrel{\mathop{:}}=1.5r_{1}^{-1}+1.5r_{2}^{-1}+r_{0}^{-1}$, and $\phi_{3}^{-1}\mathrel{\mathop{:}}=1.5r_{1}^{-1}+1.5r_{3}^{-1}+r_{0}^{-1}$, let $\phi_{\min}\mathrel{\mathop{:}}=\min(\phi_{1},\phi_{2},\phi_{3})$.
Assumption 3.
(i). $[\ln(p)]^{(2/\phi_{min})-1}=o(T)$, and (ii). $K^{2}=o(T)$, (iii). $K=o(p)$.
Assumption 4.
@@
(i). $Eigmin[\boldsymbol{c}ov(\boldsymbol{f}_{t})]\geq c>0$, with $\boldsymbol{c}ov(\boldsymbol{f}_{t})$ being the covariance matrix of the factors $\boldsymbol{f}_{t}$, $t=1,\cdots,T$. (ii). $\max_{1\leq k\leq K}\mathbb{E}\left[f_{kt}^{2}\right]\leq C<\infty$, $\min_{1\leq k\leq k}\mathbb{E}\left[f_{kt}^{2}\right]\geq c>0$. (iii). $\max_{1\leq j\leq p}\mathbb{E}\left[\eta_{j,t}^{2}\right]\leq C<\infty$.
Assumption 5.
(i).
$\max_{1\leq j\leq p}\max_{1\leq k\leq K}|b_{jk}|\leq C<\infty.$
(ii).
$\|p^{-1}\boldsymbol{B}^{\prime}\boldsymbol{B}-\boldsymbol{\Phi}\|_{l_{2}}=o(1),$
@@
for some $K\times K$ symmetric positive definite matrix $\boldsymbol{\Phi}$ such that $Eigmin(\boldsymbol{\Phi})$ is bounded away from zero.
Assumption 6.
@@
Define the rate
$l_{T}\mathrel{\mathop{:}}=r_{T}^{2}K^{5/2}\max(\bar{s}\lambda,\bar{s}^{1/2}K^{1/2}\sqrt{\frac{\max[lnp,lnT]}{T}}).$ (2)
@@
Given that $\boldsymbol{\mu}\mathrel{\mathop{:}}=(\mu_{1},\cdots,\mu_{j},\cdots,\mu_{p})^{\prime}$ is the expected portfolio return vector and $\boldsymbol{m}$ being the benchmark portfolio, and $\boldsymbol{w}_{d}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}-\boldsymbol{m}$, with $\boldsymbol{w}$ being the portfolio of interest
Assumption 7.
(i). $\bar{s}^{3}l_{T}r_{T}^{2}K^{8}=o(1)$.
@@
(ii). $\frac{\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}}{p}\geq c>0,\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\geq c>0$, with $c$ being a positive constant.
(iii). $\|\boldsymbol{m}\|_{1}=O(\|\boldsymbol{w}_{d}^{*}\|_{1})$, $\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}\geq c>0$, $|\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}|\geq c>0$.
@@
Assumptions 1-4 are used in Caner et al. (2023) to provide the consistency of the precision matrix estimator for errors. Note that Assumptions 1-3 are standard assumptions and are also used in Fan et al. (2011) as well. Also, we get $10<\phi_{2}<1,0<\phi_{3}<$ 1 given Assumption 2(iv). Furthermore, by Assumption 3, $\sqrt{\frac{\ln(p)}{T}}=o(1)$. Note that, Stationary GARCH models with finite second moments and continuous error distributions, as well as causal ARMA processes with continuous error distributions, and a certain class of stationary Markov chains satisfy our Assumptions 1-2 and are discussed in p.61 of Chang et al. (2018).
Assumption 4(i)-(ii) is also used in Fan et al. (2011), and the nodewise error assumption 4(iii) is used in Caner and Kock (2018). Next, we provide assumptions that are used to obtain consistency of the estimate of the precision matrix of returns in Caner et al. (2023). Assumption 5 is used in Fan et al. (2011). It puts some structure on the factor loadings, and this is also used in Caner et al. (2023). Assumption 6 allows the maximal eigenvalue of $\boldsymbol{\Sigma}_{u}$ to grow with T, and is used also Gagliardini et al. (2016), and also used in Caner et al. (2023). Assumption 7(i) is a tradeoff between sparsity, number of factors and eigenvalue conditions. This is more restrictive than sparsity Assumption 8 in Caner et al. (2023). Assumption 7(ii) shows that square of the maximum Sharpe Ratio (unconstrained) is lower bounded by a positive constant. This is scaled by $p$ since the numerator is summed over $p$. Then the second part of Assumption 7(ii) imposes that Global Minimum Variance (unconstrained) portfolio variance is finite. Assumption 7(ii) is used in Caner et al. (2023). Note that (16) of Caner et al. (2023) finds $\lambda=O(\sqrt{lnp/T})$ under Assumptions 1-4. Assumption 7(iii) assigns the benchmark portfolio to be growing at most as the optimal tracking error portfolio in $l_{1}$ norm in Section 3 below, and this is just needed to show that benchmark and optimal portfolios have similar behavior in $l_{1}$ norm. The rest of Assumption 7 constrains the optimal variance to be positive, and absolute portfolio return to be bounded away from zero. To show that Assumption 7 is plausible, we can have the following example, without losing any generality, with $p=2T,\bar{s}=lnT,K=lnT,r_{T}=T^{1/9}$, then
$l_{T}=T^{2/9}(ln(T))^{5/2}\max[ln(T)\sqrt{ln(2T)/T},(ln(T))^{1/2}(ln(T))^{1/2}\sqrt{ln(2T)/T}]=T^{2/9}(ln(T))^{7/2}\sqrt{ln2T/T}.$
Then
$\bar{s}^{3}l_{T}r_{T}^{2}K^{8}=T^{4/9}(ln(T))^{11+7/2}\sqrt{ln(2T)/T}\to 0.$
We also show the portfolios that are defined in Sections 2-4, and the rates that are used. Since the notation is heavy in this paper, that may help the reader. All the definitions of mentioned portfolios and rates of convergence are given in Sections below. Section 2 the portfolios are $\boldsymbol{w},\boldsymbol{w}^{*},\boldsymbol{m},\boldsymbol{w}_{d},\boldsymbol{w}_{d}^{*}$, and the rates are $r_{T}\to\infty,T\to\infty,l_{T}\to 0,T\to\infty$, as well as $r_{w1}\to 0,T\to\infty$. We do not have subscripts for $\bar{s},K$ since they can be constants or growing with $T$. Section 3 portfolios are $\boldsymbol{l},\boldsymbol{k},\boldsymbol{a},\boldsymbol{w}_{cp}^{*},\boldsymbol{w}_{R}^{*},\boldsymbol{w}_{c}^{*}$. The rate in Section 3 is $r_{w2}\to 0,T\to\infty$. For Section 4, portfolios are $\boldsymbol{w}_{op}^{*},\boldsymbol{w}_{est}^{*}$.
We differentiated our rates of convergence of weights under different constraints as $r_{w1},r_{w2}$ and did not put an additional subscript $T$, all the other rates have $T$ subscript.
3 Tracking Error Constraint
We want to maximize the expected return with a tracking error constraint as in Bajeux-Besnainou et al. (2011).
$\max_{\boldsymbol{w}\in R^{p}}\boldsymbol{\mu}^{\prime}\boldsymbol{w}\quad\mbox{subject to}\quad(\boldsymbol{w}-\boldsymbol{m})^{\prime}\boldsymbol{\Sigma}_{y}(\boldsymbol{w}-\boldsymbol{m})\leq TE^{2},\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}=1,$

### v2

@@ 2.1 Assumptions
In this subsection, we provide the assumptions. But before that we need the following notation. We use the following notation in the paper regarding vector and matrix norms. Let $\boldsymbol{v}$ be a generic $n\times 1$ vector and $\boldsymbol{H}$ to be a generic $n\times n$ matrix. Then $\|\boldsymbol{v}\|_{1},\|\boldsymbol{v}\|_{2},\|\boldsymbol{v}\|_{\infty}$ are the $l_{1}$, Euclidean, and sup norms of the vector $\boldsymbol{v}$ respectively. Next, $\|\boldsymbol{H}\|_{l_{1}},\|\boldsymbol{H}\|_{l_{\infty}}$ are the maximum column sum matrix norm and maximum row sum matrix norm, respectively. This means $\|\boldsymbol{H}\|_{l_{1}}:=\max_{1\leq j\leq n}\sum_{i=1}^{n}|H_{i,j}|$, and $\|\boldsymbol{H}\|_{l_{\infty}}:=\max_{1\leq i\leq n}\sum_{j=1}^{n}|H_{i,j}|$, where $H_{i,j}$ represents $(i,j)$ th element of matrix $H$. Let $\|\boldsymbol{H}\|_{l_{2}}$ be the spectral norm of matrix H, and $\|\boldsymbol{H}\|_{\infty}:=\max_{1\leq i\leq n}\max_{1\leq j\leq n}|H_{i,j}|$. First, $\boldsymbol{u}_{-j,t}$ is the $(p-1)\times 1$ vector of errors in $t$ th time period, except the $j$ th term in $\boldsymbol{u}_{t}$. Then define $\eta_{j,t}:=u_{j,t}-\boldsymbol{u}_{-j,t}^{\prime}\boldsymbol{\gamma}_{j}$. Also, define $Eigmin(\boldsymbol{A})$ to be the minimum eigenvalue of the matrix $\boldsymbol{A}$}.
Assumption 1
@@
(i). $\{\boldsymbol{u}_{t}\}_{t=1}^{T},\{\boldsymbol{f}_{t}\}_{t=1}^{T}$ are sequences of (strictly) stationary and ergodic random vectors. Furthermore, $\{\boldsymbol{u}_{t}\}_{t=1}^{T},\{\boldsymbol{f}_{t}\}_{t=1}^{T}$ are independent. $\boldsymbol{u}_{t}$ is a ($p\times 1$) zero mean random vector with covariance matrix $\boldsymbol{\Sigma}_{u}$ ($p\times p$). $\textnormal{Eigmin}(\boldsymbol{\Sigma}_{u})\geq c>0$, with $c$ a positive constant, and $\max_{1\leq j\leq p}E\left[u_{j,t}^{2}\right]\leq C<\infty$. (ii). For the strong mixing vector of random variables $\boldsymbol{f}_{t},\boldsymbol{u}_{t}$: $\alpha(t)\leq\exp(-Ct^{r_{0}})$, for a positive constant $r_{0}>0$.
Assumption 2
@@
There exists positive constants $r_{1}$, $r_{2}$, $r_{3}>0$ and another set of positive constants $c_{1}$, $c_{2}$, $c_{3}$, $s_{1}$, $s_{2}$, $s_{3}>0$, and for $t=1,\cdots,T$, and $j=1,\cdots,p$, with $k=1,\cdots,K$, (i). $P\left[|u_{j,t}|>s_{1}\right]\leq\exp[-(s_{1}/c_{1})^{r_{1}}].$ (ii). $P\left[|\eta_{j,t}|>s_{2}\right]\leq\exp[-(s_{2}/c_{2})^{r_{2}}].$ (iii). $P\left[|f_{k,t}|>s_{3}\right]\leq\exp[-(s_{3}/c_{3})^{r_{3}}].$ (iv). There exists $0<\phi_{1}<1$ such that $\phi_{1}^{-1}=3r_{1}^{-1}+r_{0}^{-1}$, and we also assume $3r_{2}^{-1}+r_{0}^{-1}>1$, and $3r_{3}^{-1}+r_{0}^{-1}>1$.
Define $\phi_{2}^{-1}:=1.5r_{1}^{-1}+1.5r_{2}^{-1}+r_{0}^{-1}$, and $\phi_{3}^{-1}:=1.5r_{1}^{-1}+1.5r_{3}^{-1}+r_{0}^{-1}$, let $\phi_{\min}:=\min(\phi_{1},\phi_{2},\phi_{3})$.
Assumption 3
(i). $\ln(p)]^{(2/\phi_{min})-1}=o(T)$, and (ii). $K^{2}=o(T)$, (iii). $K=o(p)$.
Assumption 4
@@
(i). $Eigmin[\boldsymbol{\Sigma}_{f}]\geq c>0$, with $\boldsymbol{\Sigma}_{f}$ being the covariance matrix of the factors $\boldsymbol{f}_{t}$, $t=1,\cdots,T$. (ii). $\max_{1\leq k\leq K}E\left[f_{kt}^{2}\right]\leq C<\infty$, $\min_{1\leq k\leq k}E\left[f_{kt}^{2}\right]\geq c>0$. (iii). $\max_{1\leq j\leq p}E\left[\eta_{j,t}^{2}\right]\leq C<\infty$.
Assumption 5
@@
(i). $\max_{1\leq j\leq p}\max_{1\leq k\leq K}|b_{jk}|\leq C<\infty.$ (ii). $\|p^{-1}\boldsymbol{B}^{\prime}\boldsymbol{B}-\boldsymbol{\Phi}\|_{l_{2}}=o(1),$ for some $K\times K$ symmetric positive definite matrix $\boldsymbol{\Phi}$ such that $Eigmin(\boldsymbol{\Phi})$ is bounded away from zero.
Assumption 6
@@
Define the rate via (16) and (22) of Caner et al. (2023)
$l_{T}:=r_{T}^{2}K^{5/2}\max(K^{2}\bar{s}^{3/2}\frac{lnp}{T},\bar{s}\sqrt{\frac{lnp}{T}},\bar{s}^{1/2}K^{1/2}\sqrt{\frac{\max[lnp,lnT]}{T}}).$ (2)
@@
We define the expected return, the benchmark, and the portfolio of interest. This helps us understand the next assumption. Define $\boldsymbol{\mu}:=(\mu_{1},\cdots,\mu_{j},\cdots,\mu_{p})^{\prime}$ as the expected return vector and $\boldsymbol{m}$ being the benchmark, and $\boldsymbol{w}_{d}^{*}:=\boldsymbol{w}^{*}-\boldsymbol{m}$, with $\boldsymbol{w}^{*}$ being the portfolio of interest.
Assumption 7
@@
(i). $\bar{s}l_{T}K^{4}=o(1)$. (ii). $\frac{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\geq c>0$, with $c$ being a positive constant. (iii). $\|\boldsymbol{m}\|_{1}=O(\|\boldsymbol{w}_{d}^{*}\|_{1})$, $\boldsymbol{w}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}^{*}\geq c>0$, $|\boldsymbol{w}^{*^{\prime}}\boldsymbol{\mu}|\geq c>0$.
@@
Assumptions 1-4 are used in Caner et al. (2023) to provide the consistency of the precision matrix estimator for errors. Also, we get $0<\phi_{2}<1,0<\phi_{3}<1$ given Assumption 2(iv). Furthermore, by Assumption 3, $\sqrt{\frac{\ln(p)}{T}}=o(1)$. Note that stationary GARCH models with finite second moments and continuous error distributions, as well as causal ARMA processes with continuous error distributions, and a certain class of stationary Markov chains satisfy our Assumptions 1-2 and are discussed in p.61 of Chang et al. (2018). Assumption 5 is used in Fan et al. (2011). It puts some structure on the factor loadings, which is also used in Caner et al. (2023).
Next, we provide assumptions that are used to obtain consistency of the estimate of the precision matrix of returns in Caner et al. (2023). Assumption 6 allows the maximal eigenvalue of $\boldsymbol{\Sigma}_{u}$ to grow with T, and is also used in Gagliardini et al. (2016), and Caner et al. (2023). Assumption 7(i) is a tradeoff between sparsity, number of factors, and eigenvalue conditions. This is more restrictive than sparsity Assumption 8 in Caner et al. (2023).
To get more intuition, we can do the following thought experiment: if we suspect that the financial markets are very noisy, $r_{T}$ being large, and if we suspect $\bar{s}$ to be large, it will not be a very good idea to have a growing number of factors in our model since the consistency of our estimators of variance and Sharpe Ratio cannot be attained. Large negative effects of a growing number of factors can be seen in estimation errors in Theorems 1-3 below. We should note that Assumption 7(i) affects the consistency of variance and Sharpe Ratio of our estimators, but consistent weight estimation is slightly easier, as seen in Theorem 1(i) versus Theorem 1(ii)-(iii). In other words, if Assumption 7(i) is violated, then the consistency of variance and Sharpe Ratio estimation is not plausible, but consistent weight estimation is plausible even when Assumption 7(i) is violated. Theorem 1(i) estimation error rate on weights are $\bar{s}l_{T}K^{3/2}$ versus Assumption 7(i) rate of $\bar{s}l_{T}K^{4}$. Another issue is the denseness of the precision matrix; what if $\bar{s}=p$? Then we can have consistency only in case of $p<T$. To delve more into Assumption 7(i), $\bar{s}l_{T}$ rate is coming from precision matrix estimation as shown in Theorem 2 of Caner et al. (2023), and $K^{3/2}$ is multiplied with precision matrix estimation rate when we have TE constraint in our portfolio weight estimation as in Theorem 1(i) below, to get the rate of convergence as $r_{w1}:=\bar{s}l_{T}K^{3/2}$. Then, to estimate variance and Sharpe Ratio, there is additional multiplication of $K^{5/2}$ as seen in Theorem 1(ii)(iii) below, which provides the rate in Assumption 7(i).
Then Assumption 7(ii) imposes that Global Minimum Variance (unconstrained-scaled) portfolio variance is finite. Assumption 7(ii) is used in Caner et al. (2023). Assumption 7(iii) assigns the benchmark portfolio to be growing at most as the optimal TE portfolio in $l_{1}$ norm in Section 3 below, and this is just needed to show that benchmark and optimal portfolios have similar behavior in $l_{1}$ norm. The rest of Assumption 7 constrains the optimal variance to be positive and absolute portfolio return to be bounded away from zero. We can have the following example to show Assumption 7(i) is plausible. Set $\bar{s}=\bar{s}(T),K=K(T)$ which are slowly varying functions of $T$. Set $r_{T}=O(T^{1/4-d_{1}})$, with $1/4>d_{1}>0$ a positive constant and $p=aT,0<a\leq C<\infty$, then, with first assuming that $\bar{s}(T)$, grows at the same rate as $K(T)$ or less, which implies the third term is the rate in (2), $l_{T}=O(T^{1/2-2d_{1}}K(T)^{3}\bar{s}(T)^{1/2}\sqrt{\frac{lnT}{T}}),$ then Assumption 7(i) is satisfied
$\bar{s}(T)^{3/2}K(T)^{7}\frac{1}{T^{2d_{1}}}\sqrt{lnT}\to 0.$
If, in the previous example, $\bar{s}(T)$, grows faster than $K(T)$, but still a slowly varying function in $T$, this implies the second term is the rate in (2) $l_{T}=O(T^{1/2-2d_{1}}K(T)^{5/2}\bar{s}(T)\sqrt{\frac{lnT}{T}})$, then
$\bar{s}(T)^{2}K(T)^{13/2}\frac{1}{T^{2d_{1}}}\sqrt{lnT}\to 0.$
In the proof of Theorem 1(i), we use in Appendix (A.20) an increasing Sharpe Ratio with the number of assets in maximum out-sample- Sharpe Ratio portfolio, $\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}$. However, we also consider two different bounded Sharpe Ratios in Appendix B. In all these two cases in Appendix B, we use a bounded Sharpe Ratio assumption for the maximum Sharpe Ratio portfolio: $0<c\leq\boldsymbol{\mu}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}\leq C<\infty$. We discuss the differences after each theorem in Remarks.
3 TE Constraint
The optimization problem is to maximize the expected return with a TE constraint:
$\max_{\boldsymbol{w}\in R^{p}}\boldsymbol{\mu}^{\prime}\boldsymbol{w}\quad\mbox{subject to}\quad(\boldsymbol{w}-\boldsymbol{m})^{\prime}\boldsymbol{\Sigma}_{y}(\boldsymbol{w}-\boldsymbol{m})\leq\text{TE}^{2},\quad\boldsymbol{1}_{p}^{\prime}\boldsymbol{w}=1,$